How do you round 0.6 to the nearest one-eighth

Answer by richwmiller(9144)

That is the strangest problem. Why eighths?
6/10=x/8
10x=48
x=48/10
x=4.8
6/10=4.8/8
or if you want 4.8 is almost 5